Small Company purchased equipment in 2005 for $100,000. Small Company did not claim any depreciation deductions on the equipment. In 2008, Small Company sold the equipment for $85,000. In March 2014, Small Company purchased new equipment for $200,000. Small Company sold the equipment in April 2015 for $220,000. How much LTCG does Small Company report in 2015?

a.

$0

b.

$20,000

c.

$220,000

d.

$5,000
Answer #1

Ans: Option d $5000 Is Correct Answer

Explanation:

Long term capital loss in the year 2008 = 100000-85000 i=$ 15000

Long term capital gain in 2015 = 220000-200000 i=$ 20000

Amount to be reported after set off in the year 2015 =$ 20000-$15000=$5000

All other Answers were Incorrect.,,

Note; As per IRS ,Capital loss can be set off for later years.